OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2). Thanks in advance.

OpenStudy (anonymous):

use tis.. (y-y1)/(x-x1)=(y2-y1)/(x2-x1) here (x1,y1)=(3,-4) (x2,y2)=(-1,2)

OpenStudy (anonymous):

Ok thank you. :)

OpenStudy (anonymous):

I got -2/-4, but I feel like thats wrong.

Nnesha (nnesha):

yes that is wrong

Nnesha (nnesha):

 use this formula